OpenStudy (anonymous):

What volume of water is required to make 0.2 M solution from 16 ml of a 0.5 M solution ?

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

\(\sf M_1V_1=M_2V_2\) => 0.2*V = 16 * 0.5 => V = 40 ml

OpenStudy (abhisar):

\(\therefore\) Volume of water needed = 40 - 16 = 24 mL

OpenStudy (anonymous):

oh! thanks...